OpenStudy (anonymous):

-2(2r + 10) = 4(9r - 5) please help and show work i don't understand:/

OpenStudy (anonymous):

When you are solving for a single variable (r) your main goal is to get it on one side of the equation. (1) decide which side you want to get r on (2) I like to circle the variables (3) look right around the circle and apply the opposite operation to get the term on the other side of the equation Where opposite operations + / - , Division / Multiplication

OpenStudy (anonymous):

Ok so first distribute and multiply:\[-4r -20 =36r -20\] Then subtract 36r from both sides:\[-40r-20=-20\] Add 20 to both sides: \[-40r = 0\] divide by -40: \[r=0\]

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

Simplifying -2(2r + 10) = 4(9r + -5) Reorder the terms: -2(10 + 2r) = 4(9r + -5) (10 * -2 + 2r * -2) = 4(9r + -5) (-20 + -4r) = 4(9r + -5) Reorder the terms: -20 + -4r = 4(-5 + 9r) -20 + -4r = (-5 * 4 + 9r * 4) -20 + -4r = (-20 + 36r) Add '20' to each side of the equation. -20 + 20 + -4r = -20 + 20 + 36r Combine like terms: -20 + 20 = 0 0 + -4r = -20 + 20 + 36r -4r = -20 + 20 + 36r Combine like terms: -20 + 20 = 0 -4r = 0 + 36r -4r = 36r Solving -4r = 36r Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Add '-36r' to each side of the equation. -4r + -36r = 36r + -36r Combine like terms: -4r + -36r = -40r -40r = 36r + -36r Combine like terms: 36r + -36r = 0 -40r = 0 Divide each side by '-40'. r = 0.0 Simplifying r = 0.0

OpenStudy (anonymous):

Hope this is all right :D